At the end of the current period, Maltese, Inc. reports the following amounts: Assets = $46,000; Liabilities = $30,000; Dividends = $2,030; Revenues = $23,600; Expenses = $11,100.

Calculate net income and stockholders' equity at the end of the period.
Answer #1
Net income = Revenues - Expenses = 23600 - 11100 =12500
Stockholders' equity = Assets - Liabilities =46000 - 30000 =16000